Calculate the present value of the following annuity streams: a. $4,000 received each year for 4 years on the last day of each year if your investments pay 5 percent compounded annually. b. $4,000 received each quarter for 4 years on the last day of each quarter if your investments pay 5 percent compounded quarterly. c.$4,000 received each year for 4 years on the first day of each year if your investments pay 5 percent compounded annually. d. $4,000 received each quarter for 4 years on the first day of each quarter if your investments pay 5 percent compounded quarterly For all requirements, do not round intermediate calculations. Round your answers to 2 decimal places. (e.g., 32.16)) a. Present value b. Present value c. Present value d. Present value 14,183.80 57,681.17 58,429.77

show on TVM solver on calculator
Answer #1

Part 1:

FV = 0

I = 5%

N = 4

PMT = 4,000

Using Financial Calculator:

PV = 14,183.80

Part 2:

FV = 0

I = 5%/ 4 = 1.25%

N = 4 * 4 = 16

PMT = 4,000

Using Financial Calculator:

PV = 57,681.17

Part 2:

FV = 0

I = 5%/ 4 = 1.25%

N = 4 * 4 = 16

PMT = 4,000

Using Financial Calculator:

PV = 57,681.17

Part 3:

Change the calcualtor to begin mode, by pressing 2nd PMT. We will see END on the screen. Press 2nd ENTER to change that to BGN and finally press 2nd CPT to exit from setting the calculation mode.

FV = 0

I = 5%

N = 4

PMT = 4,000

Using Financial Calculator:

PV = 14,892.99

Part 4:

FV = 0

I = 5%/ 4 = 1.25%

N = 4 * 4 = 16

PMT = 4,000

Using Financial Calculator:

PV = 58,402.18